Consumer Application: Televisions and computer screens are usually advertised based on the length of their diagonals. If the height of a computer screen is 11in. and the width is 14in., What is the length of the diagonal? Round to the nearest inch.
Found 2 solutions by rfer, solver91311:
Answer by rfer(10432)

c^2+a^2+b^2
c^2=11^2+14^2
c^2=121+196
c^2=317
c=18 in

Answer by solver91311(12126)



Presuming the screen is a rectangle and therefore has right angle corners, Pythagoras says:



You can run a calculator as well as I can.

John